OpenStudy (anna_spencer):

Paige borrows $4000 at a rate of 12% interest per year. What is the amount due at the end of 3 years if the interest is compounded continuously?

OpenStudy (betterthanvalentyn):

Do you have a formula to do this?

OpenStudy (anna_spencer):

No, Sorry that's why i can't figure it out.

OpenStudy (anna_spencer):

OK, Thank You!

OpenStudy (betterthanvalentyn):

tell me the answer when you finished working it out :)

OpenStudy (anna_spencer):

Well I got '5,440', but I don't know if that's it.

OpenStudy (betterthanvalentyn):

okay, i got something different. Type what you put into your calculator on here please

OpenStudy (betterthanvalentyn):

oh and i got 5619.712 btw

OpenStudy (anna_spencer):

Paige borrows $4000 at a rate of 12% interest per year. What is the amount due at the end of 3 years if the interest is compounded continuously? $4,480 $5733.31 $5,733.32 $5,440 Well i'm on Odysseyware and this is the question and the only answers

OpenStudy (betterthanvalentyn):

right, i have no idea then xD sorry i cant help you :(

OpenStudy (betterthanvalentyn):

@breanna1999 can you help Anna please

OpenStudy (breanna1999):

The formula for compounded continuously is A=Pe^rt where p= initial amount r= rate t= time

OpenStudy (breanna1999):

A=4000e^0.12(3) A=4000e^0.36

OpenStudy (breanna1999):

A=4000(1.43333) A=5,733.32